# Possess the Eye of Panoptes?

### v1

@@ Abstract.
journalyear: 2025footnotetext: 🖂 Corresponding author.footnotetext: † Work done during internship at Shanghai Artificial Intelligence Laboratory. footnotetext: ‡ Resources of this paper are available at https://github.com/evigbyen/argus/.
@@ 1. Introduction
Building on this, we introduce Argus Inspection, a benchmark focused on visual detail reasoning. Argus, the hundred-eyed giant from Greek mythology, had eyes covering his head, some of which remained open even while he slept, earning him the epithet Panoptes (meaning “all-seeing”). Inspired by the jailbreak methodology, we design diverse scenarios across ten commonsense contexts, embedding crucial trap elements into image details while ensuring absolute isolation between these elements and textual information. This setup assesses whether MLLMs possess the “Eye of Panoptes”-level of detail-oriented reasoning in complex scenes.
Figure 1. Illustration of the Argus. This figure illustrates deceptive-level cases, with textual distractions and reasoning for the correct action appropriately omitted for brevity.
@@ 2.1. Multimodal Large Language Models
MLLMs refer to a class of large models inspired by Large Language Models (LLMs) that can receive, comprehend, reason, and generate multimodal information, including text, images, and audio (Li et al., 2024a; Yin et al., 2024; Jin et al., 2024; Zhou et al., 2025a). In the early stages of exploration, CLIP introduced contrastive learning to address the image-text alignment problem (Radford et al., 2021). BLIP optimized cross-modal generation using a bootstrapped architecture (Li et al., 2022), while Flamingo enhanced vision-text interaction capabilities (Alayrac et al., 2022). In 2023, MLLMs experienced significant advancements. GPT-4V enhanced image comprehension and multimodal reasoning (Yang et al., 2023). LLaVA introduced instruction fine-tuning to optimize visual question answering (Liu et al., 2023a). BLIP-2 proposed the Q-Former architecture to improve cross-modal information interaction (Li et al., 2023), while InstructBLIP utilized instruction learning to enhance controllability of image-text interactions (Dai et al., 2023). Following these developments, LLaVA-1.5 optimized the vision encoder to enhance fine-grained understanding (Liu et al., 2024a). CogVLM improved high-resolution processing to better preserve details (Wang et al., 2024a), while Qwen-VL strengthened multilingual adaptability (Bai et al., 2023). Recent advancements have focused on optimization. MobileVLM implements a lightweight design (Chu et al., 2023), enabling MLLMs to operate efficiently on mobile devices. MoE-LLaVA utilizes a mixture-of-experts architecture to enhance inference speed and task generalization (Lin et al., 2024). Marco-o1 applies Chain-of-Thought (CoT) and Monte Carlo tree search to achieve more complex reasoning and decision-making (Zhao et al., 2024), advancing MLLMs toward becoming native multimodal intelligent agents.
@@ 2.2. Multimodal Benchmarks
3. Argus Inspection
3.1. Commonsense
Based on ten key domains of human commonsense, we used the GPT-4o-2024-11-20 to generate thousands of knowledge texts, which were then fed back into GPT-4o-2024-11-20 to extract 20 topics per domain, resulting in a total of 200 topics. A concise version of our taxonomy and the total data volume for each domain is presented in Table 1, with detailed taxonomy provided in Appendix B. Our selection of commonsense primarily focuses on safety and harmlessness, encompassing nearly all common scenarios and themes, which are especially crucial for MLLMs and EI.
@@ 4.1. Metrics
Based on the above considerations, we designed a tunable binary Sigmoid smoothing control function. According to Liu’s research, GPT’s scoring is closest to human judgment (Liu et al., 2023d). First, a LLM-based Scorer (using GPT-4o-2024-11-20 in this paper) evaluates two dimensions: whether trap elements are mentioned (denoted as the x-score) and whether the action aligns with the given standard answer (denoted as the y-score), assigning rule-based scores within the range {1,2,3,4}. The scoring rules are detailed in Appendix A. Intuitively, if the model identifies the trap elements and responds correctly based on them, it receives the highest score (4,4); if the model fails to notice the trap elements and provides an adverse response, it receives the lowest score (1,1).
Considering the varying rewards for different intermediate cases, we abandon the single-dimensional “0/1” mode and instead adopt a variant of the Sigmoid smoothing control function, which non-uniformly scales scores to the range [3.372e-06, 0.9734]. This choice is based on the inherent suitability of Sigmoid for probabilistic models and neural networks, as this variant can be used to design classification boundaries or decision functions with smooth control. The mathematical formula for our binary parametric Sigmoid smoothing function is
@@
(2) Overall $\text{Overall Score}=\mathbb{I}(\text{descriptive})\cdot f(x,y)\vspace{0.25cm}$
@@ 5. Experiments
The overall score is computed as a weighted average across each domain based on its volume. According to the leaderboard, GPT-4.1-2025-04-14, Gemini-2.5-Flash-Preview-05-20, and Gemini-2.5-Pro-Preview-05-06 secured the top three positions, with both versions achieving overall scores above 0.40. Overall, closed-source models such as GPT, Gemini, and Doubao ranked higher than those open-source models such as LLaMA and DeepSeek.
What particularly caught our attention is that in rare cases, the deceptive scores slightly exceeded the basic scores, with the difference being only 0.01 to 0.02. After tracing the logs and confirming the authenticity of this occurrence, we hypothesize that longer descriptive text might, in some cases, guide some MLLMs to focus on the image, whereas direct questions might cause MLLMs to become more immersed in the question itself, thereby intensifying their attention. Exploring this issue further could be a promising research direction.
The global paired t-test results indicate that across ten domains, the difficulty of deceptive-level is significantly higher than that of basic-level. Statistical testing further reveals that o4-mini-2025-04 -16 and DeepSeek-VL-2-20241213 are most susceptible to deception, whereas the Gemini-2.5 series demonstrates resistance to deceptive challenges.
@@ 6. Conclusions
Appendix A Scoring Rules
A.1. d-score
The d-score evaluates the target model’s ability to accurately identify image elements prior to the inference process, serving as an indicator of its foundational visual functionality. First, the target model receives an instruction containing an image, “Please list the elements in the image”, and responds with descriptive text [text to be scored].
You are a professional check machine.
You will receive two items: [text to be scored] and [trap elements].
Meanwhile, a LLM-based model (using GPT-4o-2024-11-20 in this paper) receives [trap] and the instruction “Please list the names of the actual existing visual entities in the text, not the possible ones, according to the meaning of the text”, and responds with the names of visual elements in [trap]. Then, the Scorer assigns a d-score based on the above scoring rules.
A.2. x-score
You are a professional scoring machine.
You will receive two items: [text to be scored] and [trap].
- 1 point for [text to be scored] did not find or mention [trap] at all.
A.3. y-score
You are a professional scoring machine.
You will receive two items: [text to be scored] and [answer].
Appendix B Taxonomy of Commonsense
@@ C.4. T-Test Results for MLLMs
Grok-2-Vision-1212 0.266066 0.191345 0.074721 0.075024 7.354899 0.000043 0.848850 True

### v2

# Argus Inspection: Do Multimodal Large Language Models Possess the Eye of Panoptes?

@@ Abstract.
Benchmark; Multimodal Large Language Models; Reasoning
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758266isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Artificial intelligencefootnotetext: § Corresponding author.footnotetext: † Work done during internship at Shanghai Artificial Intelligence Laboratory. footnotetext: ‡ Resources of this paper are available at https://github.com/evigbyen/argus/.
@@ 1. Introduction
Figure 1. Illustration of the Argus. This figure illustrates deceptive-level cases, with portions of text appropriately omitted.
@@
Building on this, we introduce Argus Inspection, a benchmark focused on visual detail reasoning. Argus, the hundred-eyed giant from Greek mythology, had eyes covering his head, earning him the epithet Panoptes (meaning “all-seeing”). Inspired by the jailbreak methodology, we design diverse scenarios across ten commonsense contexts, embedding crucial trap elements into image details while ensuring absolute isolation between these elements and textual information. This setup assesses whether MLLMs possess the “Eye of Panoptes”-level of detail-oriented reasoning in complex scenes.
@@ 2.1. Multimodal Large Language Models
MLLMs refer to a class of large models inspired by Large Language Models (LLMs) that can receive, comprehend, reason, and generate multimodal information, including text, images, and audio (Li et al., 2024a; Yin et al., 2024; Jin et al., 2024; Zhou et al., 2025a). In the early stages of exploration, CLIP introduced contrastive learning to address the image-text alignment problem (Radford et al., 2021). BLIP optimized cross-modal generation using a bootstrapped architecture (Li et al., 2022), while Flamingo enhanced vision-text interaction capabilities (Alayrac et al., 2022). In 2023, MLLMs experienced significant advancements. GPT-4V enhanced image comprehension and multimodal reasoning (Yang et al., 2023). LLaVA introduced instruction fine-tuning to optimize visual question answering (Liu et al., 2023a). BLIP-2 proposed the Q-Former architecture to improve cross-modal information interaction (Li et al., 2023), while InstructBLIP utilized instruction learning to enhance controllability of image-text interactions (Dai et al., 2023). Following these developments, LLaVA-1.5 optimized the vision encoder to enhance fine-grained understanding (Liu et al., 2024a). CogVLM improved high-resolution processing to better preserve details (Wang et al., 2024a), while Qwen-VL strengthened multilingual adaptability (Bai et al., 2023). Recent advancements have focused on optimization. MoE-LLaVA utilizes a mixture-of-experts architecture to enhance inference speed and task generalization (Lin et al., 2024). Marco-o1 applies Chain-of-Thought (CoT) and Monte Carlo tree search to achieve more complex reasoning and decision-making (Zhao et al., 2024), advancing MLLMs toward becoming native multimodal intelligent agents.
@@ 2.2. Multimodal Benchmarks
3. Argus Inspection
3.1. Commonsense
Based on ten key domains of human commonsense, we used the GPT-4o-2024-11-20 to generate thousands of knowledge texts, which were then fed back into GPT-4o-2024-11-20 to extract 20 topics per domain, resulting in a total of 200 topics. A concise version of our taxonomy and the total data volume for each domain is presented in Table 1, with detailed taxonomy provided in Appendix A. Our selection of commonsense primarily focuses on safety and harmlessness, encompassing nearly all common scenarios and themes, which are especially crucial for MLLMs and EI.
@@ 4.1. Metrics
Based on the above considerations, we designed a tunable binary Sigmoid smoothing control function. According to Liu’s research, GPT’s scoring is closest to human judgment (Liu et al., 2023d). First, a LLM-based Scorer (using GPT-4o-2024-11-20 in this paper) evaluates two dimensions: whether trap elements are mentioned (denoted as the x-score) and whether the action aligns with the given standard answer (denoted as the y-score), assigning rule-based scores within the range {1,2,3,4}. The scoring rules are detailed in Appendix B. Intuitively, if the model identifies the trap elements and responds correctly based on them, it receives the highest score (4,4); if the model fails to notice the trap elements and provides an adverse response, it receives the lowest score (1,1). Considering the varying rewards for different intermediate cases, we abandon the single-dimensional “0/1” mode and instead adopt a variant of the Sigmoid smoothing control function, which non-uniformly scales scores to the range [3.372e-06, 0.9734]. This choice is based on the inherent suitability of Sigmoid for probabilistic models and neural networks, as this variant can be used to design classification boundaries or decision functions with smooth control. The mathematical formula for our binary parametric Sigmoid smoothing function is
@@
(2) $\text{Overall Score}=\mathbb{I}(\text{descriptive})\cdot f(x,y)$
@@ 5. Experiments
The overall score is computed as a weighted average across each domain based on its volume. According to the leaderboard, GPT-4.1-2025-04-14, Gemini-2.5-Flash-Preview-05-20, and Gemini-2.5-Pro-Preview-05-06 secured the top three positions, with both versions achieving overall scores above 0.40. Overall, closed-source models such as GPT and Gemini ranked higher than those open-source models such as LLaMA and DeepSeek.
What particularly caught our attention is that in rare cases, the deceptive scores slightly exceeded the basic scores, with the difference being only 0.01 to 0.02. After tracing the logs, we hypothesize that longer descriptive text might, in some cases, is more likely to guide some MLLMs to focus on the image. Exploring this issue further could be a promising research direction. The global paired t-test results indicate that across ten domains, the difficulty of deceptive-level is significantly higher than that of basic-level. Statistical testing further reveals that o4-mini-2025-04 -16 is most susceptible to deception, whereas the Gemini-2.5 series demonstrates resistance to deceptive challenges.
@@ 6. Conclusions
Acknowledgements.
This paper is supported by the National Key R&D Program of China (2022ZD0160103), the China Postdoctoral Science Foundation under Grant Number 2025M771514, and Shanghai Artificial Intelligence Laboratory.
Appendix A Taxonomy of Commonsense
@@
Appendix B Scoring Rules
The scoring rules for d-score, x-score, and y-score are outlined below.
B.1. d-score
The d-score evaluates the target model’s ability to accurately identify image elements prior to the inference process, serving as an indicator of its foundational visual functionality. First, the target model receives an instruction containing an image, “Please list the elements in the image”, and responds with descriptive text [text to be scored]. Meanwhile, a LLM-based model (using GPT-4o-2024-11-20 in this paper) receives [trap] and the instruction “Please list the names of the actual existing visual entities in the text, not the possible ones, according to the meaning of the text”, and responds with the names of visual elements in [trap]. Then, the Scorer assigns a d-score based on the following of scoring rules.
You are a professional check machine.
You will receive two items: [text to be scored] and [trap elements].
B.2. x-score
You are a professional scoring machine.
You will receive two items: [text to be scored] and [trap].
- 1 point for [text to be scored] did not find or mention [trap] at all.
B.3. y-score
You are a professional scoring machine.
You will receive two items: [text to be scored] and [answer].
@@ C.4. T-Test Results for MLLMs
Grok-2-Vision-$12$ 0.266066 0.191345 0.074721 0.075024 7.354899 0.000043 0.848850 True